# The Structure of Hypergraphs Arising in Cellular Mobile Communication Systems

### v3

@@ 5 Non-Realizable Hypergraphs: Preliminaries
Recall that the worst-case performance of the maximal scheduling algorithm under the hypergraph interference model is characterized by the interference degree of the hypergraph. In the special case the hypergraph $(V,\mathcal{E})$ is $2$-uniform, its interference degree is equal to the interference degree of the graph $(V,\mathcal{E})$ [18, p. 2955]. Thus, the interference degree of the $2$-uniform hypergraph $K_{1,r}$ is $r$. In light of these results, when investigating performance guarantees of the distributed maximal scheduling algorithm for hypergraph interference models, a question that arises naturally is: what is the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable? This question is investigated in subsequent sections for various values of the path loss exponent $\alpha$. In this section, we first show that hypergraph models can be viewed as a generalization of the unit disk graph model (cf. Lemma 8). We will then prove some preliminary results that hold for any value of the path loss exponent $\alpha$ and that will be used in the proofs in subsequent sections. In the rest of this section, the path loss exponent $\alpha$ can be assumed to be any value greater than or equal to $1$.
Definition 1 of the hypergraph of a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ can be generalized by restricting the maximum size of a hyperedge to be at most some value $k$. In the special case $k=|S|$, there is no restriction on the size of a hyperedge, and so we get Definition 1 above. In the special case $k=2$, the resulting hypergraph is a unit disk graph, and conversely (see Lemma 8 below) every unit disk graph can be obtained as the hypergraph of a wireless network with $k=2$; thus, we can view the unit disk graph construction as a special case of a more general method of construction.
@@
Suppose $\mathcal{A}=\langle S,\alpha,\beta\rangle$ is a wireless network, where $\alpha\geq 1$ and $\beta=1$. Let $G=(S,E)$ be the unit disk graph generated by $S$, and let $H=(S,\mathcal{E})$ be the hypergraph generated by $\mathcal{A}$. Then, $\mathcal{E}\cap S^{(2)}=E$.
Proof: Let $\{x,y\}\subseteq S$. The Euclidean distance $d(x,y)\leq 1$ if and only if $\frac{1}{d(x,y)^{\alpha}}\geq 1$, which is the case if and only if $\{x,y\}$ is a forbidden set. Every $2$-subset of $S$ that is forbidden is a minimal forbidden set and belongs to $\mathcal{E}$. Thus, $\{x,y\}\in E$ if and only if $\{x,y\}\in\mathcal{E}$. Thus, $E=\mathcal{E}\cap S^{(2)}$ and $E\subseteq\mathcal{E}$.
As a corollary, we obtain the following: Given any unit disk graph $G=(S,E)$, there exists a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ such that the hypergraph $H=(S,\mathcal{E})$ generated by $\mathcal{A}$ satisfies $\mathcal{E}\cap S^{(2)}=E$. Conversely, given any wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ and the hypergraph $H=(S,\mathcal{E})$ generated by $\mathcal{A}$, the graph $G=(S,\mathcal{E}\cap S^{(2)})$ is a unit disk graph.
Let $\mathcal{A}=\langle S,\alpha,\beta\rangle$ be a wireless network and let $\pi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2},s\mapsto s^{\prime}$ be any map that preserves distances. For example, $\pi$ could be a rigid motion such as a rotation, a reflection, or a translation. Let $S^{\prime}:=\{\pi(s):s\in S\}$ be the set of new locations of the stations. Because $\pi$ preserves distances, it is clear that the hypergraph generated by $\mathcal{A}=\langle S,\alpha,\beta\rangle$ and the hypergraph generated by $\mathcal{A}^{\prime}=\langle S^{\prime},\alpha,\beta\rangle$ are isomorphic.
@@
Given a wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ and a positive real number $\rho$, define the scaled wireless network $\mathcal{A}^{\prime}:=\langle S^{\prime},\alpha,\beta^{\prime}\rangle$, where $S^{\prime}=\{s^{\prime}:s\in S\}$, $s^{\prime}=T_{\rho}(s)$ is the new location of station $s$ obtained using the scaling transformation $T_{\rho}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ that maps $s=(x,y)$ to $s^{\prime}:=(\rho x,\rho y)$, and $\beta^{\prime}:=\frac{\beta}{\rho^{\alpha}}$. Then, the hypergraphs generated by $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are isomorphic.
Proof: See [21].
In the rest of this section, we prove some results that are used in the proofs in subsequent sections. The reader may skip the rest of this section now and return it it later.
@@
Proof: See [21].
@@
Let $H$ be a hypergraph. If $H$ is realizable, then $H$ is realizable for some wireless network $\mathcal{A}=\langle S,\alpha,\beta\rangle$ having $\beta=1$.
Proof: See [21].
6 Nonrealizability of the Hypergraph $K_{1,5}$ When $\alpha=4$
The structure of hypergraphs generated by wireless networks $\langle S,\alpha,\beta\rangle$ depends on the value of the path loss exponent $\alpha$. In typical wireless environments, the path loss exponent $\alpha$ takes values in the range from about $1.6$ to $6.5$ [23]. The special case $\alpha=4$ has been studied by various authors [39] [42]. In the present section, we recall results from our previous work on the structure of realizable and nonrealizable hypergraphs under the signal propagation model $\alpha=4$. We first show that the $2$-uniform hypergraph $K_{1,4}$ is realizable. Then, we prove that the $2$-uniform hypergraph $K_{1,5}$ is not realizable. In other words, it can be shown that for any placement of six stations $S=\{s_{0},s_{1},\ldots,s_{5}\}\subseteq\mathbb{R}^{2}$ in the Euclidean plane and for any reception threshold $\beta$, the hypergraph generated by the wireless network $\langle S,\alpha=4,\beta\rangle$ is not isomorphic to $K_{1,5}$. Thus, while the maximal value of $r$ such that the graph $K_{1,r}$ is a unit disk graph is $r=5$, the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=4$.
@@ 6 Nonrealizability of the Hypergraph $K_{1,5}$ When $\alpha=4$
If the path loss exponent $\alpha=4$, then the hypergraph $K_{1,4}$ is realizable.
Proof: See [21].
@@
If the path loss exponent $\alpha=4$, then the hypergraph $K_{1,5}$ is not realizable.
Proof: See [21].
@@
The largest value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=4$.
@@
Before delving further, let us provide a concise overview of the proof methodology employed in [21] to establish Theorem 14, as this approach will be recurrently employed in subsequent sections. The proof of Theorem 14 unfolds in two distinct segments. In the initial segment, a proof by contradiction is constructed, detailed in [21, Section 5.2], where the assumption of the realizability of $K_{1,5}$ is made. The subsequent arguments establish that, consequently, there exist three stations on the unit circle, such that the interference at the middle station resulting from transmissions by its two closest neighbors, is deemed tolerable. In other words, there exists a value $\delta_{1}$ in the domain of the function $f(\delta_{1})$ quantifying this interference such that $f(\delta_{1})$ is less than $1$; see [21, Section 5.2] for the definition of this function. The second segment of the proof, delineated in [21, Proof of Lemma 14], demonstrates that the minimum value of $f(\delta_{1})$ over the domain exceeds $1$. This two-part proof technique will be subsequently employed to establish the nonrealizability of star hypergraphs $K_{1,r}$ for varying values of $\alpha$. Although the initial segment of the proof remains unchanged, mutatis mutandis, the function derived depends on the parameters $\alpha$ and $r$, leading to a departure from the previous scenario. Consequently, the second segments of the proofs given in the sequel are novel, necessitating the determination of the minimum value of the function for these new parameters.
7 Nonrealizability of the Hypergraph $K_{1,5}$ When $\alpha=3$
In this section, we consider the case where the path loss exponent $\alpha=3$. It will be shown that the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=4$.

### v4

@@ 5 Non-Realizable Hypergraphs: Preliminaries
In the special case the hypergraph $(V,\mathcal{E})$ is $2$-uniform, its interference degree is equal to the interference degree of the graph $(V,\mathcal{E})$ [18, p. 2955]. Thus, the interference degree of the $2$-uniform hypergraph $K_{1,r}$ is $r$. In light of these results, when investigating performance guarantees of the distributed maximal scheduling algorithm for hypergraph interference models, a question that arises naturally is: what is the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable? This question is investigated in subsequent sections for various values of the path loss exponent $\gamma$. In this section, we first show that hypergraph models can be viewed as a generalization of the unit disk graph model (cf. Lemma 10). We will then prove some preliminary results that hold for any value of the path loss exponent $\gamma$ and that will be used in the proofs in subsequent sections. In the rest of this section, the path loss exponent $\gamma$ can be assumed to be any value greater than or equal to $1$.
Definition 1 of the hypergraph of a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ can be generalized by restricting the maximum size of a hyperedge to be at most some value $k$. In the special case $k=|S|$, there is no restriction on the size of a hyperedge, and so we get Definition 1 above. In the special case $k=2$, the resulting hypergraph is a unit disk graph, and conversely (see Lemma 10 below) every unit disk graph can be obtained as the hypergraph of a wireless network with $k=2$; thus, we can view the unit disk graph construction as a special case of a more general method of construction.
@@
Suppose $\mathcal{A}=\langle S,\gamma,\beta\rangle$ is a wireless network, where $\gamma>0$ and $\beta=1$. Let $G=(S,E)$ be the unit disk graph generated by $S$, and let $H=(S,\mathcal{E})$ be the hypergraph generated by $\mathcal{A}$. Then, $\mathcal{E}\cap S^{(2)}=E$.
Proof: Let $\{x,y\}\subseteq S$. The Euclidean distance $d(x,y)\leq 1$ if and only if $\frac{1}{d(x,y)^{\gamma}}\geq 1$, which is the case if and only if $\{x,y\}$ is a forbidden set. Every $2$-subset of $S$ that is forbidden is a minimal forbidden set and belongs to $\mathcal{E}$. Thus, $\{x,y\}\in E$ if and only if $\{x,y\}\in\mathcal{E}$. Thus, $E=\mathcal{E}\cap S^{(2)}$ and $E\subseteq\mathcal{E}$.
As a corollary, we obtain the following: Given any unit disk graph $G=(S,E)$, there exists a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ such that the hypergraph $H=(S,\mathcal{E})$ generated by $\mathcal{A}$ satisfies $\mathcal{E}\cap S^{(2)}=E$. Conversely, given any wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ and the hypergraph $H=(S,\mathcal{E})$ generated by $\mathcal{A}$, the graph $G=(S,\mathcal{E}\cap S^{(2)})$ is a unit disk graph.
Let $\mathcal{A}=\langle S,\gamma,\beta\rangle$ be a wireless network and let $\pi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2},s\mapsto s^{\prime}$ be any map that preserves distances. For example, $\pi$ could be a rigid motion such as a rotation, a reflection, or a translation. Let $S^{\prime}:=\{\pi(s):s\in S\}$ be the set of new locations of the stations. Because $\pi$ preserves distances, it is clear that the hypergraph generated by $\mathcal{A}=\langle S,\gamma,\beta\rangle$ and the hypergraph generated by $\mathcal{A}^{\prime}=\langle S^{\prime},\gamma,\beta\rangle$ are isomorphic.
@@
Given a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ and a positive real number $\rho$, define the scaled wireless network $\mathcal{A}^{\prime}:=\langle S^{\prime},\gamma,\beta^{\prime}\rangle$, where $S^{\prime}=\{s^{\prime}:s\in S\}$, $s^{\prime}=T_{\rho}(s)$ is the new location of station $s$ obtained using the scaling transformation $T_{\rho}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ that maps $s=(x,y)$ to $s^{\prime}:=(\rho x,\rho y)$, and $\beta^{\prime}:=\frac{\beta}{\rho^{\gamma}}$. Then, the hypergraphs generated by the wireless networks $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are isomorphic.
Proof: Let $W\subseteq S$, and let $W^{\prime}\subseteq S^{\prime}$ be the image of $S$ under the scaling transformation $T_{\rho}$. It suffices to show that $W$ is a forbidden set for the wireless network $\mathcal{A}$ if and only if $W^{\prime}$ is a forbidden set for the wireless network $\mathcal{A}^{\prime}$, because that would imply the map $f:S\rightarrow S^{\prime},s\mapsto s^{\prime}$ is an isomorphism between the corresponding hypergraphs.
Fix $w\in W$. Then, the energy at $w^{\prime}$ due to transmissions by the set of stations $W^{\prime}-\{w^{\prime}\}$ is
$E(W^{\prime}-\{w^{\prime}\},w^{\prime})$ $=\sum_{s^{\prime}\in W^{\prime}-\{w^{\prime}\}}\frac{1}{d(s^{\prime},w^{\prime})^{\gamma}}$
$=\sum_{s\in W-\{w\}}\frac{1}{(\rho d(s,w))^{\gamma}}$
$=\frac{1}{\rho^{\gamma}}E(W-\{w\},w).$
Thus, $W^{\prime}$ is a forbidden subset for the wireless network $\mathcal{A}^{\prime}$ if and only if $W$ is a forbidden subset for the wireless network $\mathcal{A}$. The map from $S$ to $S^{\prime}$ that takes $s$ to $s^{\prime}$ is an isomorphism from the hypergraph generated by $\mathcal{A}$ to the hypergraph generated by $\mathcal{A}^{\prime}$.
In the rest of this section, we prove some results that are used in the proofs in subsequent sections. The reader may skip the rest of this section now and return to it later.
@@
Proof: Let $a,b$ be two points in a $60^{\circ}$ sector of the unit ball, with polar coordinates $a=(r_{1},\theta_{1})$ and $b=(r_{2},\theta_{2})$, where $0\leq r_{1},r_{2}\leq 1$, and where $30\leq\theta_{1},\theta_{2}\leq\pi/$ 3 without loss of generality. It will be shown that the Euclidean distance $d(a,b)$ is at most $1$.
If $r_{1}=0$ or $r_{2}=0$, then $d(a,b)\leq 1$ because the radius of the unit ball is $1$. So suppose $r_{1},r_{2}>0$. Let $\psi=|\theta_{1}-\theta_{2}|$. By the law of cosines,
$d(a,b)$ $=r_{1}^{2}+r_{2}^{2}-2r_{1}r_{2}\cos\psi$
$\leq r_{1}^{2}+r_{2}^{2}-2r_{1}r_{2}\cos\pi/3$
$=r_{1}^{2}+r_{2}^{2}-r_{1}r_{2}$
$=\frac{r_{1}^{3}+r_{2}^{3}}{r_{1}+r_{2}}$
$\leq 1,$
where the first inequality holds because $\psi\leq\pi/3$, and the second inequality holds because $r_{1}^{3}\leq r_{1}$ and $r_{2}^{3}\leq r_{2}$ when $0\leq r_{1},r_{2}\leq 1$. Equality holds throughout if and only if $r_{1}=1,r_{2}=1$ and $\psi=\pi/3$.
@@
Let $H$ be a hypergraph. If $H$ is realizable, then $H$ is realizable for some wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ having $\beta=1$.
Proof: Suppose the hypergraph $H$ is realizable. Then, there exists a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ that generates $H$. Define the scaling factor $\rho:=e^{\frac{\ln\beta}{\gamma}}$. Let $\mathcal{A}^{\prime}=\langle S^{\prime},\gamma,\beta^{\prime}\rangle$ be the wireless network obtained by scaling the wireless network $\mathcal{A}$ by a factor $\rho$ in the manner defined in the statement of Lemma 11. By Lemma 11, the wireless network $\mathcal{A}^{\prime}=\langle S^{\prime},\gamma,1\rangle$ generates $H$.
6 Nonrealizability of the Hypergraph $K_{1,5}$ When $\gamma=4$
The structure of hypergraphs generated by wireless networks $\langle S,\gamma,\beta\rangle$ depends on the value of the path loss exponent $\gamma$. In typical wireless environments, the path loss exponent $\gamma$ takes values in the range from about $1.6$ to $6.5$ [25]. The special case $\gamma=4$ has been studied by various authors [41] [45]. In the present section, we recall results from our previous work [21] on the structure of realizable and nonrealizable hypergraphs under the signal propagation model $\gamma=4$. We first show that the $2$-uniform hypergraph $K_{1,4}$ is realizable. Then, we prove that the $2$-uniform hypergraph $K_{1,5}$ is not realizable. In other words, it can be shown that for any placement of six stations $S=\{s_{0},s_{1},\ldots,s_{5}\}\subseteq\mathbb{R}^{2}$ in the Euclidean plane and for any reception threshold $\beta$, the hypergraph generated by the wireless network $\langle S,\gamma=4,\beta\rangle$ is not isomorphic to $K_{1,5}$. Thus, while the maximal value of $r$ such that the graph $K_{1,r}$ is a unit disk graph is $r=5$, the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=4$.
@@ 6 Nonrealizability of the Hypergraph $K_{1,5}$ When $\gamma=4$
If the path loss exponent $\gamma=4$, then the hypergraph $K_{1,4}$ is realizable.
Proof: Consider the wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$, where $S=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$, $s_{0}=(0,0)$, $s_{1}=(1,0)$, $s_{2}=(0,1)$, $s_{3}=(-1,0)$, and $s_{4}=(0,-1)$. Suppose $\gamma=4$ and $\beta=1$. Then, $\frac{1}{d(s_{0},s_{i})^{4}}\geq\beta$ $(i\in[4])$, and so $\{s_{0},s_{i}\}$ is a forbidden set for all $i\in[4]$.
We show that $W=\{s_{1},s_{2},s_{3},s_{4}\}$ is feasible. By symmetry, it suffices to show that the energy $E(W-\{s_{1}\},s_{1}\}$ at $s_{1}$ due to transmissions by $\{s_{2},s_{3},s_{4}\}$ is less than the reception threshold $\beta$. We have that
$E(W-\{s_{1}\},s_{1})$ $=\sum_{w\in\{s_{2},s_{3},s_{4}\}}\frac{1}{d(w,s_{1})^{4}}$
$=\frac{1}{(\sqrt{2})^{4}}+\frac{1}{2^{4}}+\frac{1}{(\sqrt{2})^{4}}<1.$
Since $W$ is feasible, the minimal forbidden sets of the wireless network $\mathcal{A}$ are precisely the subsets $\{s_{0},s_{i}\}$ ($i\in[4]$). Thus, the hypergraph generated by the wireless network $\mathcal{A}$ is the $2$-uniform hypergraph $K_{1,4}$.
@@
If the path loss exponent $\gamma=4$, then the hypergraph $K_{1,5}$ is not realizable.
Proof: By way of contradiction, suppose there exists a wireless network $\mathcal{A}=\langle S,\gamma,\beta\rangle$ that generates the hypergraph $(S,\mathcal{E})=K_{1,5}$, where $S=\{s_{0},s_{1},\ldots,s_{5}\}$ and $\mathcal{E}$ is the collection $\{\{s_{0},s_{i}\}:i\in[5]\}$ of five hyperedges. Recall that $\gamma=4$. By Lemma 14, it can be assumed that $\beta=1$. Since a translation of the Euclidean plane is a rigid motion and preserves distances, it can be assumed that the station $s_{0}$ is located at the origin.
$E(W-\{w\},w)=\sum_{s\in W-\{w\}}\frac{1}{d(s,w)^{4}}<1.$
We will arrive at a contradiction by proving that this inequality is violated for every set of five locations in the unit ball (minus the origin) for $W$.
The points in $W$ can be pushed radially outward to the boundary of the unit ball, i.e. to the unit circle, and by Lemma 13 this transformation does not decrease the distance between any two points in $W$ and hence preserves feasibility of $W$. Thus, we can assume that the hypergraph $K_{1,5}$ is generated by a wireless network $\langle S,\gamma,\beta\rangle$, such that $S=\{s_{0},s_{1},\ldots,s_{5}\}$, $\gamma=4$, $\beta=1$, and the set of points $W=\{s_{1},\ldots,s_{5}\}$ lies on the unit circle.
Figure 1: Five stations on the unit circle, with stations $s_{1},s_{2}$ and $s_{5}$ close enough to each other that $\delta_{1}+\delta_{5}=2\pi/15$. Relevant to the proof of Theorem 16.
$f(\delta_{1})=\frac{1}{d(s_{1},s_{2})^{4}}+\frac{1}{d(s_{1},s_{5})^{4}},$
where the domain for $f$ is $[0,2\pi/15]$. By Lemma 17, the minimum value of $f(\delta_{1})$ is greater than $1$. Because the reception threshold $\beta$ is equal to $1$, the subset $\{s_{1},s_{2},s_{5}\}$ is a forbidden subset. Hence, its superset $W$ is also a forbidden set, contradicting the fact that $W$ is an independent set.
Lemma 17.
Let $f(\delta_{1})$ be the function
$f(\delta_{1}):=\frac{1}{d(s_{1},s_{2})^{4}}+\frac{1}{d(s_{1},s_{5})^{4}},$
Define the function $g(\delta)=d(s_{1},s_{2})^{-4}+d(s_{1},s_{5})^{-4}$, where $s_{1}=(1,0)$, $s_{2}=(1,2\pi/5+\delta)$, $s_{5}=(1,2\pi/5-\delta)$, and the domain is $[-\pi/15,\pi/15]$. Because $f(\delta_{1})=g(\delta+\pi/15)$, it suffices to show that the minimum value of $g(\delta)$ is greater than $1$.
It can be verified that
$g(\delta)=\frac{1}{4\left(1-\cos\left(\frac{2\pi}{5}+\delta\right)\right)^{2}}+\frac{1}{4\left(1-\cos\left(\frac{2\pi}{5}-\delta\right)\right)^{2}},$
$g^{\prime}(\delta)=-\frac{\sin\left(\frac{2\pi}{5}+\delta\right)}{2\left(1-\cos\left(\frac{2\pi}{5}+\delta\right)\right)^{3}}+\frac{\sin\left(\frac{2\pi}{5}-\delta\right)}{2\left(1-\cos\left(\frac{2\pi}{5}-\delta\right)\right)^{3}},$
and
$g^{\prime\prime}(\delta)=\frac{3-\cos\left(\frac{2\pi}{5}+\delta\right)-2\cos^{2}\left(\frac{2\pi}{5}+\delta\right)}{2\left(1-\cos\left(\frac{2\pi}{5}+\delta\right)\right)^{4}}$
$+\frac{3-\cos\left(\frac{2\pi}{5}-\delta\right)-2\cos^{2}\left(\frac{2\pi}{5}-\delta\right)}{2\left(1-\cos\left(\frac{2\pi}{5}-\delta\right)\right)^{4}}.$
Because $03-\cos\gamma-2\cos^{2}\gamma>$ 0 when $|\cos\gamma|<1$, $g^{\prime\prime}(\delta)>0$ for all $\delta$. Also, $g^{\prime}(0)=0$. Thus, $g(\delta)$ is convex and attains a global minimum at $\delta=0$. It can be verified that $g(0)=\frac{3+\sqrt{5}}{5}$, which is greater than $1$.
@@
If the path loss exponent $\gamma=4$, then the largest value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=4$.
@@
Combining Theorem 16 and Proposition 28, we obtain a generalization of Theorem 16 to all nonintegral values of the path loss exponent $\gamma\leq 4$: if the path loss exponent $\gamma\leq 4$, then the hypergraph $K_{1,5}$ is not realizable.
7 Nonrealizability of the Hypergraph $K_{1,5}$ When $\gamma=3$
In this section, we consider the case where the path loss exponent $\gamma=3$. It will be shown that the maximal value of $r$ such that the hypergraph $K_{1,r}$ is realizable is $r=4$.